# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ 2.3. Reinforcement Learning
Reinforcement learning is a framework for addressing the problem of a RL-Agent learning to interact with an environment, formalized as a Markov Decision Process (MDP) $\mathcal{M}:=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of possible actions, $\mathcal{T}$ is the transition function, $\mathcal{R}(s,a)$ is the reward function, and $\gamma\in[0,1]$ is a discount factor that determines the importance of future rewards. In this framework, the RL-Agent and the environment interact at discrete time steps. At the time step $t$, the RL-Agent observes a state $s_{t}\in\mathcal{S}$ from the environment. Subsequently, the RL-Agent selects an action $a\in\mathcal{A}$ based on a policy $\pi$ which is defined as a probability distribution over actions for each state, i.e., $\pi(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ given the state $s_{t}$. Upon taking action $a_{t}$, the RL-Agent receives a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$ and transitions to a new state $s_{t+1}$ according to the transition function $\mathcal{T}$. The policy $\pi$ is updated based on the observed interactions between the RL-Agent and the environment.
There are several methods for updating the policy. In the policy-based reinforcement learning, the policy is parameterized and updated directly using classical optimizers (i.e. gradient descent). On the other hand, the value-based reinforcement learning, instead of directly learning the policy $\pi$, learns to estimate the state-value function which is calculated as the expected sum of discounted future rewards given a policy $\pi$. Subsequently, the policy for action selection is updated in a greedy manner based on the state-value function. The state-value function is expressed as follows:
(3) $\mathcal{V}^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$
Following this approach, it is common to use the action-value function or Q-function, denoted as $\mathcal{Q}^{\pi}(s,a)$, which represents the expected cumulative reward if the RL-Agent chooses action $a$ under the current state $s$ and the policy $\pi$. Given the state $s^{\prime}$ obtained by taking the action $a$ under the state $s$, the Q-function is presented as follows:
(4) $\mathcal{Q}^{\pi}(s,a)=E_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s,a_{t}=a\right]=\mathcal{R}(s,a)+\gamma\mathcal{V}^{\pi}(s^{\prime})$
Recently, hybrid approaches leverage the advantages of both policy-based and value-based techniques by concurrently learning both a policy and a state-value function. One prominent method in this approach is the Actor-Critic method which is essential for addressing the exploration-exploitation trade-off problem in reinforcement learning. In the Actor-Critic paradigm, two key models are employed: the Actor, parameterized by $\theta_{A}$, which predicts the policy $\pi\left(a\mid s\right)$, and the Critic, parameterized by $\theta_{C}$, which predicts the state value function $\mathcal{V}^{\pi}(s)$. The Policy Gradient algorithm is commonly utilized to update the parameters $\theta_{A}$ and $\theta_{C}$. The objective function employed in the Policy Gradient method is expressed as follows:
@@
(5) $L^{PG}(\theta_{A},\theta_{C})=\hat{\mathbb{E}}_{t}\left[\log\pi\left(a_{t}\mid s_{t}\right)(\mathcal{Q}^{\pi}(s_{t},a_{t})-\mathcal{V}^{\pi}(s_{t}))\right]$
@@ 3. Proposed Solutions
In this section, we initially present our baseline RL framework, called NaiveRL, and CHARME, our advanced solution (Sect. 3.1). The three key components of CHARME are introduced in subsequent sections 3.2-3.4 along with its theoretical analysis.
@@ 3.1. RL framework
In this part, we first present NaiveRL which is an initial RL framework for the minor embedding problem. In this framework, solutions are constructed by sequentially mapping one node from a given logical graph to one node in a hardware graph. This approach gives an initial intuition of how to formulate an optimization problem by an RL framework. We then identify drawbacks of this framework which corresponds to challenges mentioned before. Accordingly, we introduce CHARME, a chain-based RL framework which effectively addresses the above drawbacks of NaiveRL. Unlike NaiveRL, CHARME  constructs solutions by sequentially mapping one node from a given logical graph to a chain of nodes in a hardware graph. By doing this, CHARME is able to maintain the satisfaction of three embedding constraints after each step of embedding, thereby ensuring the feasibility of the resulting solutions.
@@ 3.1.1. NaiveRL  - An Initial RL framework
We consider a simple RL framework which constructs solutions by sequentially embedding one node in logical graph $P$ to one unembedded node in the hardware graph $H$. The algorithm terminates when the found solution satisfies three embedding constraints: chain connection, global connection, and one-to-many connection. Alternatively, it stops if there is no remaining unembedded node in $H$. We define the states, actions, transition function and reward function in the RL framework as follows:
@@
States: A state $s$ is a combination of the logical graph $P$, the hardware graph $H$, and a embedding $\phi:V_{P}\rightarrow\mathcal{P}(V_{H})$. Specifically, we denote a state at step $t$ as $s_{t}=\{P,H,\phi^{(t)}\}$.
@@
Actions: An action is in form of $(u,v)$ with $u\in V_{H}$ and $v\in V_{P}$. Note that $v$ is an unembedded node which is not in current embedding $\phi$. The action $(u,v)$ is equivalent to embedding node $v$ in $P$ to an unembedded node $u$ in $H$.
@@
Transition function: When an action $(u,v)$ is taken, a new solution $\phi^{\prime}$ with $\phi^{\prime}(v)=\phi(v)\cup\{u\}$ is obtained. The new state is a combination of logical graph $P$, hardware graph $H$, and the new solution $\phi^{\prime}$.
@@
Reward function: At a step $t$, the reward $r_{t}$ for selecting an action $(u,v)$ consists of two terms. The first term, named Verification $\vartheta_{t}$, is determined based on the quality of resulting solution after taking an action $(u,v)$. Specifically, if the solution $\phi^{(t)}$ satisfies all three embedding constraints, $\vartheta_{t}$ is a number of actions taken to reach that state (which is equivalent to $t$). Otherwise, $\vartheta_{t}$ is set to 0.
@@
In addition, to address the challenge of sparse rewards caused by the large search space, we introduce an additional component for the reward function, named Exploration $\varepsilon_{t}$. This component is designed to encourage the RL framework to initially learn from existing solutions before independently exploring the search space. In particular, $\varepsilon_{t}$ is calculated by the difference between the action found by our NaiveRL  framework at step $t$, denoted as $a_{t}$, and the action found by an existing step-by-step heuristic method (Ngo et al., 2023a) at step $t$, denoted as $\bar{a}_{t}$. The extent to which exploration contributes to the reward function is controlled by a variable $\sigma\in[0,1]$. In the initial stages of the training phase, we set the variable $\sigma$ as 0. It implies that in the early steps of training, the policy of the RL framework mimics the heuristic strategy. As the policy is progressively updated with precomputed solutions, $\sigma$ gradually decreases in proportion to the agent’s success rate in finding solutions that satisfy the three embedding constraints. To sum up, the reward function $r_{t}$ for a given state $s_{t}$ can be expressed as follows:
(6) $r_{t}=-\vartheta_{t}-\varepsilon_{t}=-\vartheta_{t}-(1-\sigma)\left(a_{t}-\bar{a}_{t}\right)^{2}$
Limitations of NaiveRL. Here, we assess the practical performance of NaiveRL, in order to identify its limitations. First, we generate a dataset of 1000 graphs and split them into training and test sets. Our goal is to train the model on 700 Barabási-Albert (BA) graphs with $n=70$ nodes and degree $d=5$ with the aim of enabling effective generalization to the remaining 300 graphs with various sizes ($n\in\{30,50\}$) and a lower degree ($d=2$). Figure 2 illustrates the limitation in the generalization ability of NaiveRL. Without the exploration term in Equation 6, the model faced difficulties in finding feasible solutions. Despite an improvement achieved by adding the exploration term in Equation 6, the model is still struggling to generalize effectively to unseen graphs. Furthermore, there is no guarantee that NaiveRL  can consistently generate feasible solutions, even with further training. This presents a bottleneck in the quantum annealing process, as feasible solutions are always necessary. As a result, there is a need for an enhanced and efficient RL framework design that can generalizes its policy for unseen graphs and ensure the feasibility of resulting solutions.
@@ 3.1.2. CHARME: A chain-based RL solution for the minor embedding problem
Figure 3. The workflow of CHARME - achain-based reinforcement learning framework
@@
Figure 4. A GNN-based architecture of the models (actor and critic) presenting the policy of the RL-Agent
@@
To overcome the limitations of NaiveRL, we introduce our proposed CHARME, a chain-based RL framework that embeds one node in the logical graph into a chain of nodes in the hardware graph at each step. CHARME  guarantees to find feasible solutions within exact $|V_{P}|$ steps. Consequently, compared to the naive approach, the chain-based RL framework significantly reduces and simplifies the search space. In the following, we describe the chain-based RL framework.
@@
States: We use the same definition of states in the former framework. Here, we specify components of states in more details. The state $s_{t}$ at step $t$ contains:
@@
  • The logical graph $P$: We represent the logical graph $P$ as a tuple $(X_{P}^{(a)},X_{P}^{(f)})$. $X_{P}^{(a)}\in\{0,1\}^{|V_{P}|\times|V_{P}|}$ indicates the adjacency matrix of $P$. Given the number of node features as $d_{P}$, $X_{P}^{(f)}\in\mathbb{R}^{|V_{P}|\times d_{P}}$ indicates the node feature matrix of $P$. We specify the node features of $P$ later.
@@
  • The hardware graph $H$: We represent the logical graph $H$ as a tuple $(X_{H}^{(a)},X_{H}^{(f)})$. $X_{H}^{(a)}\in\{0,1\}^{|V_{H}|\times|V_{H}|}$ indicates the adjacency matrix of $H$. Given the number of node features as $d_{H}$, $X_{H}^{(f)}\in\mathbb{R}^{|V_{H}|\times d_{H}}$ indicates the node feature matrix of $H$. We specify the node features of $H$ later.
@@
  • The solution/embedding: We denote $\phi^{(t)}:V_{P}\rightarrow\mathcal{P}(V_{H})$ as an embedding from $P$ to $H$ at step $t$. Specifically, if a node $v\in V_{P}$ is embedded onto a node $u\in V_{H}$, then $u\in\phi(v)$.
@@
Actions: At step $t$, we denote $U^{(t)}_{P}$ as a set of embedded node in $P$. In each step, In each step, an action $a_{t}$ is defined as choosing an unembedded node $v\in V_{P}\setminus U^{(t)}_{P}$. From here, for simplicity, we consider each action as an unembedded node of $P$. Based on $a_{t}$, we can find the next feasible embedding $\phi^{(t+1)}$ of $P[U^{(t)}_{P}\cup\{a_{t}\}]$ in $H$.
@@
Transition function: Because the given graphs $P$ and $H$ remain unchanged throughout an episode (which is explained later), we consider a transition from state $s^{(t)}$ to state $s^{(t+1)}$ as the transition from $\phi^{(t)}$ to $\phi^{(t+1)}$. The algorithm for state transition will be discussed in the subsequent section.
@@
Reward function: At step $t$, a reward $r_{t}$ for selecting a node $u\in P$ is determined based on two components. The first component is the qubit gain compared to the current embedding, denoted as $c_{t}$. This component can be computed as the difference between the number of qubits of $\phi^{(t+1)}$ and $\phi^{(t)}$. The second component is the exploration term which is similarly defined as that in the reward function of NaiveRL, with the difference being that the precomputed solution $\bar{a}_{t}$ is obtained through our proposed exploration algorithm, which we will discuss later. This component plays a role in facilitating the training process of the chain-based RL framework. To sum up, the reward $r_{t}$ is defined as follows:
(7) $r_{t}=-c_{t}-\varepsilon_{t}=-(|\phi^{(t+1)}|-|\phi^{(t)}|)-(1-\sigma)\left(a_{t}-\bar{a}_{t}\right)^{2}$
@@
In order to describe the workflow of CHARME, first of all, we denote a round of finding a solution for a pair of $(P,H)$ as an episode. Given a training set of logical and hardware graphs, CHARME  runs multiple episodes to explore solutions of pairs of logical and hardware graphs from the training set. In one episode, the framework constructs a solution through a series of steps. Figure 3 shows the workflow of CHARME  in a step $t$. Specifically, at step $t$ of an episode, CHARME  receives the state $s_{t}$ including the logical graph, the hardware graph and the corresponding embedding, as inputs. The RL-Agent uses its policy to map the given state to an appropriate action $a_{t}$, leading to the next state (Note that the policy of the RL-Agent is also the policy of the RL framework). In our CHARME, we apply the Actor-Critic method (Mnih et al., 2016) for our policy. The actor and the critic are represented by two different models with a similar structure which is explained later. Next, the state transition algorithm takes current state $s_{t}$ and chosen action $a_{t}$ as inputs to produces the next state $s_{t+1}$. Finally the environment updates state $s_{t}$ to $s_{t+1}$, and calculates the reward $r_{t}$. Information in step $t$ including states, actions, and rewards is stored in a roll-out buffer which is later used for updating the RL-Agent’s policy. The episode continues until the RL-Agent reaches the terminal state where no unembedded node remains.
@@
It is observable that for each episode, the RL-Agent explores an embedding order $O=\{a_{1},\dots,a_{|V_{P}|}\}$ for a training pair $(P,H)$. By learning from reward signals along with embedding orders, CHARME  is able to generate efficient embedding orders for unseen logical graphs and hardware graphs, consequently reducing the qubit usage.

### v2

@@ 2.3. Reinforcement Learning
Reinforcement learning (RL) is a framework for addressing the problem of an RL-Agent learning to interact with an environment, formalized as a Markov Decision Process (MDP) $\mathcal{M}:=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of possible actions, $\mathcal{T}$ is the transition function, $\mathcal{R}(s,a)$ is the reward function, and $\gamma\in[0,1]$ is a discount factor that determines the importance of future rewards. In this framework, the RL-Agent and the environment interact in discrete time steps. At time step $t$, the RL-Agent observes a state $s_{t}\in\mathcal{S}$ from the environment. Subsequently, the RL-Agent selects an action $a\in\mathcal{A}$ based on a policy $\pi:\mathcal{S}\rightarrow\mathbb{P}(\mathcal{A})$ which defines a probability distribution over actions given each state. Formally, the action is sampled as $a_{t}\sim\pi(\cdot\mid s_{t})$. Upon executing action $a_{t}$, the RL agent receives a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$ and transitions to a new state $s_{t+1}$ according to the transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$. The tuple of $(s_{t},a_{t},r_{t},s_{t+1})$ is then stored in a roll-out buffer $\mathcal{D}$, which is used to update policy $\pi$ later.
Next, we present the method for policy optimization. To begin, it is essential to define what is an optimal policy. A common metric for evaluating a policy is the state-value function, which quantifies the expected cumulative reward when following policy $\pi$ from a given state.
Formally, given a policy $\pi$, the state-value function for state $s\in\mathcal{S}$ is defined as the expected sum of future discounted rewards:
(3) $\mathcal{V}^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s\right],$
where $\gamma\in(0,1]$ is the discount factor that controls the trade-off between immediate and future rewards.
The goal of RL is to find the optimal policy $\pi^{*}$ that maximizes the long-term expected return for any initial state $s_{0}\in\mathcal{S}$. This can be formulated as the following optimization problem:
$\pi^{*}=\arg\max_{\pi}\mathcal{V}^{\pi}(s_{0}).$
Actor-Critic (Mnih et al., 2016) is a specific type of Policy Gradient methods (Sutton et al., 1999) designed to find the optimal policy. In this method, the policy is represented by a parameterized learning model actor, denoted as $\hat{\pi}_{\theta}$, where $\theta$ represents the trainable parameters. A key advantage of the Actor-Critic method over other policy gradient approaches is the inclusion of critic, another parameterized learning model that estimates the state-value function $\mathcal{V}^{\pi}$. Specifically, the critic is represented as $\hat{\mathcal{V}}_{\omega}$, where $\omega$ denotes its trainable parameters. The critic helps reduce the variance of policy gradient estimates, leading to more stable learning.
The actor model is trained by minimizing the following loss:
@@
$\mathcal{L}_{actor}(\theta)$ $=\mathbb{E}_{\hat{\pi}_{\theta},a_{t}\sim\hat{\pi}_{\theta}(\cdot|s_{t}),s_{t+1}\sim\mathcal{T}(s_{t},a_{t})}\left[\log\hat{\pi}_{\theta}\left(a_{t}\mid s_{t}\right)(\mathcal{R}(s_{t},a_{t})+\gamma\hat{\mathcal{V}}_{\omega}(s_{t+1})-\hat{\mathcal{V}}_{\omega}(s_{t}))\right]$
(4) $\approx\frac{1}{|\mathcal{D}|}\sum_{(s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{D}}\left[\log\hat{\pi}_{\theta}\left(a_{t}\mid s_{t}\right)(r_{t}+\gamma\hat{\mathcal{V}}_{\omega}(s_{t+1})-\hat{\mathcal{V}}_{\omega}(s_{t}))\right]$
On the other hand, the critic model is trained by minimizing the following loss:
$\mathcal{L}_{critic}(\omega)$ $=\mathbb{E}_{\hat{\pi}_{\theta},a_{t}\sim\hat{\pi}_{\theta}(\cdot|s_{t}),s_{t+1}\sim\mathcal{T}(s_{t},a_{t})}\left[(r_{t}+\gamma\hat{\mathcal{V}}_{\omega}(s_{t+1})-\hat{\mathcal{V}}_{\omega}(s_{t}))^{2}\right]$
(5) $\approx\frac{1}{|\mathcal{D}|}\sum_{(s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{D}}\left[(r_{t}+\gamma\hat{\mathcal{V}}_{\omega}(s_{t+1})-\hat{\mathcal{V}}_{\omega}(s_{t}))^{2}\right]$
The actor and critic are updated simultaneously. Given the update factors for the actor and critic as $\alpha_{1}$ and $\alpha_{2}$, the parameters $\theta$ of the actor model and $\omega$ of the critic model can be updated as:
(6) $\theta\leftarrow\theta+\alpha_{1}\nabla_{\theta}\mathcal{L}_{actor}(\theta)$
(7) $\omega\leftarrow\omega+\alpha_{2}\nabla_{\omega}\mathcal{L}_{critic}(\omega)$
@@ 3. Proposed Solutions
In this section, we present our baseline RL framework, NaiveRL, and our advanced solution, CHARME (Section 3.1). The three key components of CHARME are introduced in Subsections 3.2-3.4 along with its theoretical analysis.
@@ 3.1. RL Framework
Here, we present NaiveRL as an initial RL framework for the minor embedding problem. In this framework, solutions are constructed by sequentially mapping one node from a given logical graph to one node in a hardware graph. This approach gives an initial intuition of how to formulate an optimization problem using an RL framework. We then identify the drawbacks of this framework that correspond to the challenges mentioned before. Accordingly, we introduce CHARME, a chain-based RL framework that effectively addresses the above drawbacks of NaiveRL. Unlike NaiveRL, CHARME constructs solutions by sequentially mapping one node from a given logical graph to a chain of nodes in a hardware graph. By doing this, CHARME is able to maintain the satisfaction of three embedding constraints after each step of embedding, thus ensuring the feasibility of the resulting solutions.
@@ 3.1.1. NaiveRL - An Initial RL Framework
We consider a simple RL framework that constructs solutions by sequentially embedding one node in logical graph $P$ to one unembedded node in the hardware graph $H$. The algorithm terminates when the solution found satisfies three embedding constraints: chain connection, global connection, and one-to-many connection. Alternatively, it stops if there is no remaining unembedded node in $H$. We define the states, actions, transition function, and reward function in the RL framework as follows:
@@
States: A state $s$ is a combination of the logical graph $P$, the hardware graph $H$, and an embedding $\phi:V_{P}\rightarrow\mathcal{P}(V_{H})$. Specifically, we denote the state at step $t$ as $s_{t}=\{P,H,\phi^{(t)}\}$.
@@
Actions: An action is in the form of $(u,v)$ with $u\in V_{H}$ and $v\in V_{P}$. Note that $v$ is an unembedded node that is not in the current embedding $\phi$. The action $(u,v)$ is equivalent to an embedding node $v$ in $P$ to an unembedded node $u$ in $H$.
@@
Transition Function: When an action $(u,v)$ is taken, a new solution $\phi^{\prime}$ with $\phi^{\prime}(v)=\phi(v)\cup\{u\}$ is obtained. The new state is a combination of the logical graph $P$, the hardware graph $H$, and the new solution $\phi^{\prime}$.
@@
Reward Function: At a step $t$, the reward $r_{t}$ for selecting an action $(u,v)$ consists of two terms. The first term, named Verification $\vartheta_{t}$, is determined based on the quality of the resulting solution after taking an action $(u,v)$. Specifically, if the solution $\phi^{(t)}$ satisfies the three embedding constraints, $\vartheta_{t}$ is the number of actions taken to reach that state (which is equivalent to $t$). Otherwise, $\vartheta_{t}$ is set to 0.
@@
In addition, to address the challenge of sparse rewards caused by the large search space, we introduce an additional component for the reward function, named Exploration $\varepsilon_{t}$. This component is designed to encourage the RL framework to initially learn from existing solutions before independently exploring the search space. In particular, $\varepsilon_{t}$ is calculated by the difference between the action found in our NaiveRL framework at step $t$, denoted as $a_{t}$, and the action found by an existing step-by-step heuristic method (Ngo et al., 2023a) at step $t$, denoted as $\bar{a}_{t}$. The extent to which exploration contributes to the reward function is controlled by a variable $\sigma\in[0,1]$. In the initial stages of the training phase, we set the variable $\sigma$ to 0. It implies that in the early steps of training, the policy of the RL framework mimics the heuristic strategy. As the policy is progressively updated with precomputed solutions, $\sigma$ gradually increases in proportion to the agent’s success rate in finding solutions that satisfy the three embedding constraints. In sum, the reward function $r_{t}$ for a given state $s_{t}$ can be expressed as follows:
(8) $r_{t}=-\vartheta_{t}-\varepsilon_{t}=-\vartheta_{t}-\left[1-\sigma\right]_{+}\left(a_{t}-\bar{a}_{t}\right)^{2}$
However, a major drawback of this naive approach is its inability to consistently produce feasible solutions that satisfy the constraints of the minor embedding problem. As demonstrated in the Experiment section, the feasibility rate remains low even for small logical graphs. This poses a significant bottleneck for the quantum annealing process, as finding a feasible embedding may require an unpredictable number of reruns. Therefore, an enhanced RL framework, that can ensure the feasibility of its solutions, is needed.
@@ 3.1.2. CHARME: A Chain-Based RL Solution for the Minor Embedding Problem
Figure 2. Workflow of the CHARME framework. The detailed architecture of the actor and critic is described in the subsequent part.
@@
Figure 3. A GCN-based architecture of the models (actor and critic) presenting the policy of the RL-Agent
@@
To overcome the limitations of NaiveRL, we introduce CHARME, a chain-based RL framework that embeds one node in the logical graph into a chain of nodes in the hardware graph at each step. CHARME guarantees to find feasible solutions within exact $|V_{P}|$ turns of action selection and state transition (proven in Subsection 3.3). Consequently, compared to the naive approach, CHARME significantly reduces and simplifies the search space. In the following, we formally describe CHARME:
@@
States: We use the same definition of states as in the former framework. Here, we specify the components of states in more detail. The state $s_{t}$ at step $t$ contains:
@@
  • The logical graph $P$: We represent the logical graph $P$ as a tuple $(X_{P}^{(a)},X_{P}^{(f)})$. $X_{P}^{(a)}\in\{0,1\}^{|V_{P}|\times|V_{P}|}$ indicates the adjacency matrix of $P$. Given the number of node features $d^{\prime}_{P}$, $X_{P}^{(f)}\in\mathbb{R}^{|V_{P}|\times d^{\prime}_{P}}$ indicates the node feature matrix of $P$. Subsequently, we specify the node features of $P$.
@@
  • The hardware graph $H$: We represent the logical graph $H$ as a tuple $(X_{H}^{(a)},X_{H}^{(f)})$. $X_{H}^{(a)}\in\{0,1\}^{|V_{H}|\times|V_{H}|}$ indicates the adjacency matrix of $H$. Given the number of node features $d^{\prime}_{H}$, $X_{H}^{(f)}\in\mathbb{R}^{|V_{H}|\times d^{\prime}_{H}}$ indicates the node feature matrix of $H$. Subsequently, we specify the node features of $H$.
@@
  • The solution/embedding: We denote $\phi^{(t)}:V_{P}\rightarrow\mathcal{P}(V_{H})$ as an embedding from $P$ to $H$ at step $t$. Specifically, if a node $v\in V_{P}$ is embedded in a node $u\in V_{H}$, then $u\in\phi(v)$.
@@
Actions: At step $t$, we denote $U^{(t)}_{P}$ as a set of embedded nodes in $P$. In each step, an action $a_{t}$ is defined as the choice of an unembedded node $v\in V_{P}\setminus U^{(t)}_{P}$. From here, for simplicity, we consider each action as an unembedded node of $P$. Based on $a_{t}$, we can find the next feasible embedding $\phi^{(t+1)}$ of $P[U^{(t)}_{P}\cup\{a_{t}\}]$ in $H$.
@@
Transition Function: Because the given graphs $P$ and $H$ remain unchanged throughout an episode, we consider a transition from state $s^{(t)}$ to state $s^{(t+1)}$ as the transition from $\phi^{(t)}$ to $\phi^{(t+1)}$. The transition state $s^{(t)}$ to state $s^{(t+1)}$ is deterministic and does not depend on the policy. The algorithm for state transition (i.e., Algorithm 1) will be discussed in the subsequent section. In general, given an inputted state, Algorithm 1 utilizes ATOM’s subroutines (Ngo et al., 2023a) (i.e., Node Embedding and Topology Adapting described in Section 2.2) to return a deterministic state as output.
@@
Reward function: At step $t$, a reward $r_{t}$ is determined to select a node $u\in P$ based on two components. The first component is the qubit gain compared to the current embedding, denoted as $c_{t}$. This component can be computed as the difference between the number of qubits of $\phi^{(t+1)}$ and $\phi^{(t)}$. The second component is the exploration term, which is similarly defined as that in the reward function of NaiveRL, with the difference being that the precomputed solution $\bar{a}_{t}$ is obtained through our proposed exploration algorithm. This component is crucial in enhancing the training process of the CHARME framework. Specifically, at the start of the training process, the policy can be updated with high-quality embedding orders pre-computed by Algorithm 2, rather than relying on randomly explored embedding orders. We recall that the action $a_{t}$ and the precomputed solution $\bar{a}_{t}$ are presented as non-negative integers corresponding to the index of the selected node in the logical graph. Additionally, the exploration term is regulated by the variable $\sigma$, which progressively increases in proportion to the agent’s success rate in finding feasible solutions. In sum, the reward $r_{t}$ is defined as follows:
(9) $r_{t}=-c_{t}-\varepsilon_{t}=-(|\phi^{(t+1)}|-|\phi^{(t)}|)-\left[1-\sigma\right]_{+}\left(a_{t}-\bar{a}_{t}\right)^{2}$
@@
In order to describe the workflow of CHARME, we denote a round of finding a solution for a pair of $(P,H)$ as an episode. Given a training set of logical and hardware graphs, CHARME runs multiple episodes to explore solutions of pairs of logical and hardware graphs from the training set. In one episode, the framework constructs a solution through a series of steps. Figure 2 shows the workflow of CHARME in an arbitrary step $t$. Specifically, at step $t$ of an episode, CHARME receives the state $s_{t}$ including the logical graph, the hardware graph, and the corresponding embedding, as input. Then, similar to NaiveRL, the actor-critic method is employed to represent and train the policy. Specifically, the actor maps the current state $s_{t}$ to an action $a_{t}$. The detailed architecture for the actor (and critic) is described in Section 3.2. Then, the state transition algorithm (Algorithm 1) takes the current state $s_{t}$ and the chosen action $a_{t}$ as input to produce the next state $s_{t+1}$. Finally, the environment updates the state $s_{t}$ to $s_{t+1}$, and calculates the reward $r_{t}$. The information in step $t$, including states, actions, and rewards, is stored in a roll-out buffer. The episode continues until the terminal state is reached where no unembedded node remains. After a certain number of training steps, all information in the roll-out buffer is used to update the actor and critic according to Equation 6 and 7. Then, a part of the roll-out buffer is randomly deleted to make space for information in subsequent episodes (i.e., $80\%$). In addition, $\sigma$ starts at 0 and increases by $0.01$ after each update until $1-\sigma]$ reaches 0.
@@
It is observable that for each episode the RL-Agent explores an embedding order $O=(a_{1},\dots,a_{|V_{P}|})$ for a training pair $(P,H)$. By learning from reward signals along with embedding orders, CHARME is able to generate efficient embedding orders for unseen logical graphs and hardware graphs, consequently reducing the usage of qubits.